please solve the formula
Solve the formula for the specified variable.
A = 1/2 bh for b
thanks

Found 2 solutions by minerva, prabhjyot:
Answer by minerva(22)

no problem,
A = 1/2 bh for b
a=bh/2
2a=bh
2a/h=b

or if the question is
a=1/2bh (bh in teh denominator)
then..
2abh=1
b=1/2ah
i hope i was helpful


Answer by prabhjyot(165)

A = (1/2)bh
A/[(1/2)h] = b

b = A/[(1/2)h]
b=2A/h